The probabilities of a machine manufacturing 0,1,2,3 and 5 defective parts in one day are 0.75,0.17,0.04,0.025,0.01 and 0.005 respectively.Find the mean of the probability distribution.

1
Expert's answer
2022-05-20T04:16:32-0400

First of all, let us restate the formula of the mean:


"E(\xi) = \Sigma_{i=1}^n p(x_i)x_i"


Mean of the given probability distribution:


"E(\xi) = p(0)*0 + p(1)*1 + p(2)*2 + p(3)*3 + p(4) *4 +p(5)*5 ="

"= 0.17 * 1 + 0.04*2+0.025*3+0.01*4+0.005*5 ="

"=0.17+0.08+0.075+0.04+0.025=0.17+0.12+0.1="

"=0.39"